# A Framework for Quantum Simulations of Energy-Loss and Hadronization in Non-Abelian Gauge Theories: SU(2) Lattice Gauge Theory in 1+1D

### v2

@@ Appendix B Details of the $L=2$ System
We use the Lanczos algorithm [Lanczos:1950zz] to generate the ground-state energies and wavefunctions. Working with $g=1.0$ and $m_{q}=0.1$ in the $n_{Q}=0$ sector, our starting state is the SC vacuum is
@@
There are three mesonic operators (two acting on a single spatial site, $\hat{O}_{M0}^{(0)}$ and $\hat{O}_{M0}^{(1)}$, and one between sites, $\hat{O}_{M1}^{(0,1)}$), and three baryon operators with analogous actions ($\hat{O}_{B0}^{(0)}$, $\hat{O}_{B0}^{(1)}$, and $\hat{O}_{B1}^{(0,1)}$) that can be used to prepare the interacting ground state with sub-percent precision,
@@
$\hat{O}_{B1}^{(0,1)}$ $\equiv-4i\hat{I}^{4}\left(\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{Z}^{2}\hat{\sigma}^{-}\hat{\sigma}^{-}-\hat{\sigma}^{-}\hat{\sigma}^{-}\hat{Z}^{2}\hat{\sigma}^{+}\hat{\sigma}^{+}\right)\hat{I}^{2}\ .$ (83)
@@ Appendix E Circuit Construction
Figure 21: Different circuits implementing $e^{-i\frac{\theta}{2}(\hat{Y}\hat{Z}^{n}\hat{X}-\hat{X}\hat{Z}^{n}\hat{Y})}$, the building blocks for implementing the mesonic operators defined in Eq. (50). The R-box composite elements in the circuit are defined in Fig. 13.
@@
Figure 22: Circuits that implement unitary transformations from the mesonic operators. The red line in first circuit denotes a cancellation between block operators. The CNOT-depth of these circuits are: 8, 28, 60, respectively. The R-box composite elements in the circuit are defined in Fig. 13.
@@
For $\hat{O}_{B1}$, additional $\hat{Z}$s are inserted in between, implemented as CNOT operators connecting the extra qubits, to retain the fermionic nature, as shown in Fig. 24.
Figure 24: A Circuit implementing the baryon operator $e^{-i\theta\hat{O}_{B1}}$ that generates an amplitude for creating a baryon-anti-baryon pair separated by one lattice site.

### v3

@@ Appendix B Details of the $L=2$ System
We use the Lanczos algorithm [183] to generate the ground-state energies and wavefunctions. Working with $g=1.0$ and $m_{q}=0.1$ in the $n_{Q}=0$ sector, our starting state is the SC vacuum is
@@
There are three meson operators (two acting on a single spatial site, $\hat{O}_{M0}^{(0)}$ and $\hat{O}_{M0}^{(1)}$, and one between sites, $\hat{O}_{M1}^{(0,1)}$), and three baryon operators with analogous actions ($\hat{O}_{B0}^{(0)}$, $\hat{O}_{B0}^{(1)}$, and $\hat{O}_{B1}^{(0,1)}$) that can be used to prepare the interacting ground state with sub-percent precision,
@@
$\hat{O}_{B1}^{(0,1)}$ $\equiv-4i\hat{I}^{4}\left(\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{I^{2}}\hat{\sigma}^{-}\hat{\sigma}^{-}-\hat{\sigma}^{-}\hat{\sigma}^{-}\hat{I^{2}}\hat{\sigma}^{+}\hat{\sigma}^{+}\right)\hat{I}^{2}\ .$ (83)
@@ Appendix E Circuit Construction
Figure 21: Different circuits implementing $e^{-i\frac{\theta}{2}(\hat{Y}\hat{Z}^{n}\hat{X}-\hat{X}\hat{Z}^{n}\hat{Y})}$, the building blocks for implementing the meson operators defined in Eq. (50). The R-box composite elements in the circuit are defined in Fig. 13.
@@
Figure 22: Circuits that implement unitary transformations from the meson operators. The red line in first circuit denotes a cancellation between block operators. The CNOT-depth of these circuits are: 8, 28, 60, respectively. The R-box composite elements in the circuit are defined in Fig. 13.